-inch television has a rectangular screen on which the diago- nal measures 25 inches. The ratio of the height to the width of the screen is 3 to 4. Let the height equal 3x and the width equal 4x, and then use the Pythagorean theorem to solve for x. Then find the height and width of this television screen.

The length of the tv is_______inches while the width is________inches

Answer by mananth(16946)

using Pythagorean theorem
(3x)^2+(4x)^2=25^2
9x^2+16x^2=625
25x^2=625
/25
x^2=25
x=5 (positive value)
Height = 3x=15 in
width = 4x = 20 in